OpenStudy (anonymous):

If 100 is 3/4 of a cup, how much is 4/4 of a cup?

OpenStudy (lgbasallote):

first, divide 100 by 3/4

OpenStudy (lgbasallote):

what number do you get?

OpenStudy (zzr0ck3r):

(3/4)x = 100 (4/4)x = x = ?

OpenStudy (anonymous):

(3/4)x = 100 x = 133.33 = total amount

OpenStudy (anonymous):

133

OpenStudy (lgbasallote):

right. so that's 4/4 cup

OpenStudy (anonymous):

Ok thanks!

OpenStudy (lgbasallote):

welcome